OpenStudy (anonymous):

Find the x intercept and the y intercept of the graph of this equation y = 5xsquared + 6x + 1

OpenStudy (anonymous):

x=0 for y intercept and y=0 for x intercepts

OpenStudy (anonymous):

you can find y intercept when x=0 and for x intercepts y=5x^2 + 6x + 1 0=5x^2 + 6x + 1 0=(5x+1)(x+1) x=-1/5 or x=-1

OpenStudy (anonymous):

Thank you!